OpenStudy (anonymous):

Myra has a square garden with an area of 52 square feet. She wants to put a chain link fence along one complete side of her garden. Fencing is sold in whole numbers of feet. What is the LEAST amount of fencing she will have to buy? A. 26 feet B. 13 feet C. 8 feet D. 7 feet

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

since it said "LEAST", I was thinking D, am I right?

ganeshie8 (ganeshie8):

Area of square = s^2 = 52

ganeshie8 (ganeshie8):

find out the side length : s

ganeshie8 (ganeshie8):

s^2 = 52 s = ?

ganeshie8 (ganeshie8):